The following costs appear in a company's flexible budget at an activity level of 15,000 machine hours:

Indirect materials $7800

Factory rent $18,000

What would be the flexible budget amounts at an activity level of 12,000 machine hours if indirect material is a variable cost and factory rent a fixed cost and actual factory rent is $22,000?

A. $7800 for indirect labour and $14,400 for factory rent

B. $7800 for indirect labour and $18,000 for factory rent

C. $6240 for indirect labour and $14,400 for factory rent

D. $6240 for indirect labour and $18,000 for factory rent

E. $6240 for indirect labour and $22,000 for factory rent

Choose one of the following:

A) Option A

B) Option B

C) Option C

D) Option D

E) Option E
Answer #1

Variable cost Per MH =total indirect material / Machine Hours = 7800 / 15000=0.52

Indirect material 12000 hours= 12000*0.52= $ 6,240

Factory rent at 12000 MH = $ 18,000

flexible Budget cost at 12000 MachineHours

Option D) $ 6,240 for Indirect Labour and $ 18,000 For Factory Rent (Answer)